An airplane is descending in preparation for a landing at an airport. The height, h, in meters of the airplane above the ground level of the airport is a linear function of time, t, where t is the number of minutes after the plane started to descend. The plane was 3,915 meters above the ground level of the airport 2 minutes into the plane's decent and was 2,025 meters above after 9 minutes. How long will it take the plane to land at the airport after it started to descend
1
Expert's answer
2020-11-12T19:33:04-0500
"Solution"

After it started to descend, it took the plane exactly "7\ minutes" to descend from "3915\ meters" to "2025\ meters"

"Initial\ Height\ (H_0)\ =\ 3915\ m\\
Final\ Height\ (H_1)\ =\ 2025\ m\\
Displacement\ =\ \delta H=1890m\\
Change\ in\ time\ =\ \delta t\ =\ 9-2=7\ minutes\ \times\ 60\ =\ 420\ s\\"

Now, we will calculate the velocity, "\bar v"


"\bar v=\frac{\delta H}{\delta t}= \frac{1890m}{420s}=4.5ms^{-1}"

Given "\bar v=4.5ms^{-1}"

"Initial\ Height\ (H_0)\ =\ 2025\ m\\
Final\ Height\ (H_1)\ =\ 0\ m\\
Displacement\ =\ \delta H=2025\ m \\"

"4.5\ ms^{-1}\ \delta t=2025\ m\\
\delta t\ =\ \frac{2025\ m}{4.5 ms^{-1}}=450\ seconds"


Total time it took the plane to land from the time it started to descend.

Total time = "(420 + 450)s=870\ seconds"


"\therefore T=14.5 \ seconds"